*Again-Rules for rounding: If you are calculating in tablets, always round to the nearest 1/2 tablet....

  1. *Again-Rules for rounding:
    • If you are calculating in tablets, always round to the nearest 1/2 tablet.
    • When you are calculating IV' rates, round to the nearest whole number.
    • Perform rounding only the end of the calculation
  2. CONVERSTIONS TO USE
  3. 1 grain =60 mg
  4. 15 grain=1 gram
  5. 1 pint= 2 cups=480mL
  6. 1 quart= 2 pt= 4 cups=960mL
  7. 1 oz= 30 mL
  8. 1 cup = 8 oz = 240 mL
  9. 1 quart = 2 pt = 4cups= 960 mL
  10. 1cc= 1mL
  11. 1tsp= 5mL
  12. 1 Tbs= 15 mL
  13. 1 gallon= 4 qt = 8 pt
  14. Note:Some drug conversions will use cc(cubic centimeters) This is the same as a mL. 1cc=1mL

(*There are actually 29.57 mL in 1 fl oz, but 30 mL is usually used by nurses in medical facilities. We will stay with this standard for all liquid measurements. 1 oz=30mL, 1 c=240mL, 1 pt=480mL, 1 quart=960mL.)

QUESTIONS:

1. The patient weighs 50 lb. The physician orders 30 mg/kg/day of a certain medication. How many mg do you give in one day?________mg/day(round to the nearest whole number)

(Remember not to round till the very end)

2. The nurse is preparing potassium bicarbonate (TriLyte) for a pediatric client scheduled for a colonoscopy. The directions state to provide the prep at an estimated rate of 25 mL/kg/hr. The child weighs 55 lb. The nurse should encourage the child to drink approximately how much of the solution each hour? (round to the nearest 1/2 a cup.)

Group of answer choices

a. 2 cups

b. 3 cups

c. 3.5 cups

d. 2.5 cups

1 cup

3. The nurse receives orders to start metronidazole (Flagyl) 50 mg/kg/24 hours divided into three equal doses for a child that weighs 33 lbs. The medication is provided in 250-mg tablets. How many tablets will the nurse provide with each dose?_____tablets
Answer #1

Question 1

Given

Weight:50lbs(0.454×50=22.7kg)

Order:30mg/kg/day

That is 1kg=30mg

Then 22.7kg=??mg

Cross multiply and divide

=22.7×30÷1

=681mg/day

Upon rounding off to nearest whole number it will be 600mg

Ans:600mg/day

Question 2

Given

Med:Potassium bicarbonate

Weight:55lbs (0.454×55=24.97kg or 25kg)

25ml/kg/hr

That is 1kg=25ml

So 25kg=??ml

Cross multiply and divide

=25×25÷1

=625ml/hr

As 1cup = 240ml

Then 625ml=??cup

Cross multiply and divide

=625×1÷240

=625÷240

=2.6 cup

Upon rounding off to the nearest half cup it is 2.5 cups

Ans:2.5 cups

3.Given

Medication:Flagyl

Order:50mg/kg/24hour

Frequency:Three times a day

Weight:33lvs(0.454×33=14.9 or 15kg)

As 1kg=50mg

Then 15kg=??mg

Cross multiply and divide

=15×50÷1

=750÷1

=750mg/24 hour or 750mg in three divided doses for a day

Available:250mg/tab

When 1tablet=250mg

Then 750mg=??tablet

Cross multiply and divide

=750×1÷250

=750÷250

=3 tablet /24 hr

So to calculate 1 time dose 750 ÷3 =250 mg ,which means 1 tablet per dose

Ans:The nurse has to provide 1 tablet in each dose